20220513_DX_Night. LBTI AO, Day 1

07:00 (UT) change configuration
  • cp bad_actuators.list.mmc bad_actuators.list
  • run idl in the directory $ADOPT_ROOT/conf/adsec/current/TS3/ and at the prompt, execute: adsec_make_bad_act_files (new files are created act_wo_pos and act_wo_curr)
  • cp act_w_mmc_20220510.txt act_w_mmc.txt
  • + 50e-9 force +0.11
  • cp -r data.20220421.telescope/* data
  • rm default.sav
  • ln -s local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav default.sav

07:34 (UT) applied flat: 631 last mode applied, actuator #356 has a slightly off force of +0.04

11:46 (UT) RB temperature 7.0 Celsius

-- GuidoBrusa - 13 May 2022
Topic revision: r3 - 13 May 2022, GuidoBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback